A bonds credit rating provides a guide to it's price. assume aaa bonds yield 3.7% and baa bonds yield 4.7%. assume a 10% five-year bond with annual coupons and a face value of $1,000. (do not round intermediate calculations. round you're answers to 2 decimal places.)
Answer #1

Calculate the present value of bond when the interest rate is 3.7% as follows:

X in f =PV(3.7%,5,-1000*10%,-1000) ? X Function Arguments Rate 3.7% Nper 51 Pmt -1000*10% Fv -1000 Type 5 * = 0.037 = 5 = -10

Present value of the bond is $1,282.84

---------------------------------------------------------------------------

Calculate the present value of bond as follows:

X in fo =PV(4.7%,5,-1000*10%,-1000) х Function Arguments PV Rate 4.7% * * Nper Pmt -1000*10% Fv -1000 Type = 0.047 = 5 = -100

Present value of the bond is $1,231.38.

--------------------------------------------------------

As the interest rate is increased value of the bond is decreased. Therefore, relationship between the bond price and interest rate is inverse.